A car starts from rest and reaches a speed of 45km/hr in 20s. it travels at this speed for Β½ a kilometre and then comes to rest uniformly in 30 s Find;

a)Acceleration

b)total distance traveles


1
Expert's answer
2022-06-01T12:49:41-0400

Given:

"v_0=0, \quad v_1=45\:\rm km/hr=12.5\: m/s"

"t_1=20\:\rm s"

"d_2=500\:\rm m"

"t_3=30\:\rm s"

"v_4=0"

(a)

"a_1=\frac{\Delta v}{\Delta t}=\frac{v_1-v_0}{t_1}=\frac{12.5-0}{20}=0.625\:\rm m/s^2"

"a_3=\frac{\Delta v}{\Delta t}=\frac{v_4-v_1}{t_3}=\frac{0-12.5}{30}=-0.417\:\rm m/s^2"

(b)

"d=d_1+d_2+d_3\\
=\frac{v_1+v_0}{2}t_1+d_2+\frac{v_4+v_1}{2}t_3\\
=\frac{12.5+0}{2}*20+500+\frac{0+12.5}{2}30=813\:\rm m"